OpenStudy (anonymous):

Make this pattern block, red, green, blue, red, green, blue. Suppose their are 20 blocks in the pattern. How many blocks would be blue? How would you know what the number of blue blocks would be?

OpenStudy (dean.shyy):

This is a ratio question. Think of the ratio like this: R:G:B:R:G:B = 2R:2G;2B = 1:1:1 However, with 20 blocks, the equality of each color is not complete. You do not get a whole number. So, if you look at the mapping, it would be: R:G:B:R:G:B = 6 block colors R:G:B:R:G:B = Another 6 block colors = 12 block colors total R:G:B:R:G:B = Another 6 block colors = 18 block colors total Now, there are two blocks to fill the 20 block count. All you need to do is to take the first two color patterns, which is R and G, which is RED and GREEN. To conclude the answer, you add up all the color blocks listed plus the last two filler color blocks. So, you would put: 2R + 2G + 2B 2R + 2G + 2B 2R + 2G + 2B 1R + 1G First, red gives you 2R + 2R + 2R + 1R = 7 REDS Second, green gives you 2G + 2G + 2G + 1G = 7 GREEN Third, blue gives you 2B + 2B + 2B = 6 BLUE That is a total of 20 blocks.

OpenStudy (anonymous):

Thank You